OpenStudy (xkryponite):

19. a box contains 95 pink rubber bands and 90 brown rubber bands. You select a rubber bands at random from the box. Find each probability. Write the probability as a fraction in simplest form. A. Find the theoretical probability of selecting a Pink rubber band. B. Find the theoretical probability of selecting a Brown rubber band. C. You repeatedly choose a rubber band from the box, record the color, and put the rubber band back in the box. The results are shown in the table below. Find the experimental probability of each color based on the table.

OpenStudy (xkryponite):

I need help with the 3rd one, my first two answers are a: 19/37 = 95/185 b: 18/37 = 90/185

OpenStudy (knov):

do you have a table given ?

OpenStudy (xkryponite):

oops sorry yeah, outcome | occurrences pink | 36 brown | 33

OpenStudy (knov):

oh, i just deleted my drawing.

OpenStudy (xkryponite):

purpose or accident xd

OpenStudy (knov):



OpenStudy (knov):



OpenStudy (xkryponite):

thanx <3